# Fast Policy Learning for 6-DOF Position Control of Underwater Vehicles

### v1

@@ Abstract
Autonomous Underwater Vehicles (AUVs) require reliable six–degree-of-freedom (6-DOF) position control to operate effectively in complex and dynamic marine environments. Traditional controllers are effective under nominal conditions but exhibit degraded performance when faced with unmodeled dynamics or environmental disturbances. Reinforcement learning (RL) provides a powerful alternative but training is typically slow and sim-to-real transfer remains challenging. This work introduces a GPU-accelerated RL training pipeline built in JAX and MuJoCo-XLA (MJX). By jointly JIT-compiling large-scale parallel physics simulation and learning updates, we achieve training times of under two minutes. Through systematic evaluation of multiple RL algorithms, we show robust 6-DOF trajectory tracking and effective disturbance rejection in real underwater experiments, with policies transferred zero-shot from simulation. Our results provide the first explicit real-world demonstration of RL-based AUV position control across all six degrees of freedom.
@@ I Introduction
  • We release the codebase to facilitate reproducibility and adoption by the community (codebase).
@@ II Related Work
A smaller set of works demonstrated real experiments. In [18], DDPG was applied for 6-DOF velocity control and successfully deployed on a physical vehicle. In [19], SAC was used for trajectory tracking, but the evaluation was restricted to translational degrees of freedom. In [20], data-informed domain randomization enabled sim-to-real transfer for translational position and yaw control. Some studies have focused on specific objectives requiring position control. For example, [21] applied SAC to a 3DoF link-alignment task, demonstrating moderate performance on a real vehicle. In [22], authors demonstrate real experiments on a station keeping task, where PPO policies are trained in IsaacLab within 15 minutes by leveraging parallelized GPU-based simulation. However, station keeping task is not fully representative of 6-DOF control capabilities. In [23], the authors present an end-to-end DRL controller for trajectory following with full 6-DOF actuation, trained over 15 hours and validated in both simulation and in-water experiments. Although the controller stabilizes roll and pitch, these axes are largely constrained to near-zero values during reported real experiments, and general roll or pitch tracking is not explored.
In summary, while RL has shown promise for AUV control, prior work either remains confined to simulation or demonstrates only partial control objectives in real settings, leaving the problem of transferring full 6-DOF control capabilities from simulation to real vehicles open.
@@ VII Conclusion
This work introduced a GPU-accelerated reinforcement learning framework for 6-DOF position control of underwater vehicles, combining differentiable simulation in MJX with large-scale parallelization and jit compilation. By jointly compiling physics and learning updates, we reduced training times up to a couple of minutes while obtaining policies that reliably achieve 6-DOF control. All RL policies demonstrated precise trajectory tracking and disturbance rejection, surpassing an MPC baseline and enabling complex manoeuvres that, to our knowledge, have not been extensively reported before. These results establish that fast policy learning and robust sim-to-real transfer are feasible for full 6-DOF underwater vehicle control. More broadly, the GPU-accelerated MJX–JAX pipeline demonstrates a general approach for fast policy learning with differentiable physics, with potential impact beyond underwater robotics.

### v2

@@ Abstract
Autonomous Underwater Vehicles (AUVs) require reliable six–degree-of-freedom (6-DOF) position control to operate effectively in complex and dynamic marine environments. Traditional controllers are effective under nominal conditions but exhibit degraded performance when faced with unmodeled dynamics or environmental disturbances. Reinforcement learning (RL) provides a powerful alternative but training is typically slow and sim-to-real transfer remains challenging. This work introduces a GPU-accelerated RL training pipeline built in JAX and MuJoCo-XLA (MJX). By jointly JIT-compiling large-scale parallel physics simulation and learning updates, we achieve training times of under two minutes. Through systematic evaluation of multiple RL algorithms, we show robust 6-DOF trajectory tracking and effective disturbance rejection in real underwater experiments, with policies transferred zero-shot from simulation.
@@ II Related Work
A smaller set of works demonstrated real experiments. In [18], DDPG was applied for 6-DOF velocity control and successfully deployed on a physical vehicle. In [19], SAC was used for trajectory tracking, but the evaluation was restricted to translational degrees of freedom. In [20], data-informed domain randomization enabled sim-to-real transfer for translational position and yaw control. Some studies have focused on specific objectives requiring position control. For example, [21] applied SAC to a 3DoF link-alignment task, demonstrating moderate performance on a real vehicle. In [22], authors demonstrate real experiments on a station keeping task, where PPO policies are trained in IsaacLab within 15 minutes by leveraging parallelized GPU-based simulation. However, station keeping task is not fully representative of 6-DOF control capabilities. In [23], the authors present an end-to-end DRL controller for trajectory following with full 6-DOF control, trained over 15 hours and validated in both simulation and in-water experiments.
In summary, RL has shown recent success in 6-DOF AUV control, however prior work mostly remains confined to simulation, and in-water validated approaches are either trained over impractical timescales or shown limited control performances.
@@ VII Conclusion
This work introduced a GPU-accelerated reinforcement learning framework for 6-DOF position control of underwater vehicles, combining differentiable simulation in MJX with large-scale parallelization and jit compilation. By jointly compiling physics and learning updates, we reduced training times up to a couple of minutes while obtaining policies that reliably achieve accurate 6-DOF control. All RL policies demonstrated precise trajectory tracking and disturbance rejection. These results establish that fast policy learning and robust sim-to-real transfer are feasible for full 6-DOF underwater vehicle control. More broadly, the GPU-accelerated MJX–JAX pipeline demonstrates a general approach for fast policy learning with differentiable physics, with potential impact beyond underwater robotics.